Show that five-fold rotational symmetry is not possible in a 2-D lattice
1
Expert's answer
2020-02-04T09:15:53-0500

Since we deal with flat 2D lattice, we need to ask a question: what regular polygons can tile the 2D plane? These figures are triangles, some quadrilaterals and hexagons. Why? Because we the interior angle of a polygon with "n" sides is


"\theta=180^\circ-(360/n)."

How to tile a 2D plane without holes? By following this condition:


"m\theta=360^\circ,"

which means that theta must divide 360 degrees into equal m integer parts. Solve these two equations and express "m(n)":


"m(n)=\frac{2n}{n-2}."

Only n equal to 3, 4 and 6 comply with this condition. That is why five-fold rotational symmetry is impossible in 2D.